Answer with chain rule


Z = cos (x² + y²), x = u cos v, y = u sin v

1
Expert's answer
2021-11-17T03:48:22-0500

"Z_u=Z_xx_u+Z_yy_u"

"Z_u=-2xsin(x\u00b2 + y\u00b2)cosv-2ysin(x\u00b2 + y\u00b2)sinv="

"=-2ucos^2vsin(u^2)-2usin^2vsin(u^2)=-2usin(u^2)"


"Z_v=Z_xx_v+Z_yy_v"

"Z_v=2xusin(x\u00b2 + y\u00b2)sinv-2yusin(x\u00b2 + y\u00b2)cosv="

"=u^2sin2vsinu^2-u^2sin2vsinu^2=0"